A selling company is a regular C corporation. Given the following data calculate the net proceeds to the shareholders of the selling firm if the buyer makes a stock acquisition versus an acquisition of assets.

      Purchase price, stock                                                        $250

      Purchase price, assets                                                     250

      Liabilities of seller                                                          100

      Basis in assets (seller)                                                       150

      Basis in shares (shareholders of seller)                             125

      Marginal corporate tax rate (federal and state)                35%

      Individual capital gains tax rate (federal and state)          24%
Answer #1

net proceeds to the shareholders of the selling firm if the buyer makes a stock acquisition:

net proceeds = Purchase price of stock - [Purchase price of stock - Basis in shares (shareholders of seller)]*[1- Individual capital gains tax rate (federal and state)]

net proceeds = $250 - ($250 - $150)*(1 - 0.24) = $250 - $100*0.76 = $250 - $76 = $174

net proceeds to the shareholders of the selling firm if the buyer makes an acquisition of assets:

net proceeds = [(Purchase price of assets - Liabilities of seller)*(1- Marginal corporate tax rate (federal and state))] - Basis in shares (shareholders of seller)

net proceeds = [($250 - $100)*(1 - 0.35)] - $125 = ($150*0.65) - $125 = $97.5 - $125 = -$27.5